# ReMamba: Equip Mamba with Effective Long-Sequence Modeling

### v2

@@ Results
We choose the English branch of LongBench because our training set only contains English. Higher values across all indicators are indicative of better performance. We compare the performance of the models in detailed tasks in Table 1 under the max length 6k corresponding to the training setting. Here the hyperparameters for ReMamba are: $s=0$, $p=0.18$ and $\rho=0.009$. We will also show later that our model’s robustness to various of hyperparameter combinations. Table 1 shows that our ReMamba model improves the average scores on LongBench 3.23 compared to the SFT Mamba baseline. Our model approaches the pretrained and finetuned transformer baseline.
@@
We compare the performance on the closed-ended tasks of L-Eval. The higher all indicators are, the better. A snap of detailed task scores for the maximum length of 6k is presented in Table 2. We can witness a 1.64 improvement on average scores compared to the SFT Mamba baseline. Here the hyperparameter setting for ReMamba is: $s=0$, $p=0.20$ and $\rho=0.05$.
Analyses and Discussions
Varying Length
To complement our main results, which employ a maximum sequence length of 6k tokens to align with training settings, we further evaluate the model performance at varying input lengths ranging from 2k to 9k tokens. This evaluation is conducted using the LongBench and L-Eval benchmarks. As depicted in Figure 3, our ReMamba consistently outperforms the baseline Mamba model across all tested context lengths on LongBench. Notably, the performance gap between our model and the baseline widens as the context length increases. Furthermore, our model extends the efficient context length (the length at which greatest performance is observed) to 6k tokens, compared to 4k tokens for the finetuned Mamba baseline. In Figure 4, we observe performance improvements across all context lengths for our model on L-Eval. Our ReMamba even surpasses the transformers baseline.
Speed Performance and Memory Expense
Our model introduces a single additional forward pass during inference, resulting in no additional memory consumption. To evaluate the speed performance, we varys the input sequence length from 1k to 8k tokens while fixing the output length at 1k tokens. For each configuration, we use a batch size of 1 and measure the speed on an NVIDIA A100 80GB GPU. We compare the performance of ReMamba, Mamba, and the vanilla transformer model (llama2-3b), as illustrated in Figure 6. The speed metric is given in tokens per second. Our experiments indicate that ReMamba operates at speeds comparable to the original baseline, maintaining a significant speed advantage over traditional transformers.
Generalizing to Mamba2
While our method is specifically tailored for Mamba, we also conduct experiments to verify its applicability to Mamba2. As is shown in Table 3, the same method applied to Mamba2 (we call ReMamba2 here) achieves 1.6 improved performance on averaged scores of LongBench. Here we use $s=0$, $p=0.25$ and $\rho=0.05$. The max length is still 6k. It is noteworthy that Mamba2 exhibits nearly no performance improvement over Mamba on LongBench, suggesting potential limitations within the Mamba model series.
Ablation Study
@@
Figure 3: Average scores on LongBench varying max length from 2k to 9k. The “Pre” means pretrained model while “SFT” means finetuned model. The performance of llama2-3b (SFT) and llama2-3b (Pre) is for reference, using the max length of 2k due to positional encoding problems.
Figure 4: Average scores on L-Eval varying max length from 2k to 9k. The performance of llama2-3b (SFT) and llama2-3b (Pre) is for reference.
@@
Robustness varying choices of hyperparamters
The aforementioned results were obtained using the hyperparameter settings $s=0$, $p=0.18$, and $\rho=0.009$, which demonstrates relatively superior performance. In Figure 7, we also show the stability of our model by varying the hyperparameters $p$ and $\rho$. For these experiments, the parameter $s$ is fixed at 0.
@@
Table 4: Performance of different model variants on LongBench. In this context, the “ReMamba” model type constitutes our optimal model. The “middle” type corresponds to the model variant where $s$ is non-zero. The “special” model variant compresses the entire prompt using $\rho=0.009$ and subsequently appends the compressed hidden states to the end of the original prompt in the second stage.
Why compress from the start

### v4

@@ 5.3 Results
We choose the English branch of LongBench because our training set only contains English. Higher values across all indicators are indicative of better performance. We compare the performance of the models in detailed tasks in Table 1 under the max length 6k corresponding to the training setting. Here the hyperparameters for ReMamba are: $s=0$, $p=0.18$ and $\rho=0.009$. We will also show later that our model’s robustness to various of hyperparameter combinations. Table 1 shows that our ReMamba model improves the average scores on LongBench 3.23 compared to the SFT Mamba baseline. Our model approaches the pretrained and finetuned transformer baseline. The results for DeciMamba indicate that it may not be sufficiently effective for tasks in LongBench or it may be sensitive to the choosing of hyperparameters.
@@
We compare the performance on the closed-ended tasks of L-Eval. The higher all indicators are, the better. A snap of detailed task scores for the maximum length of 6k is presented in Table 2. We can witness a 1.64 improvement on average scores compared to the SFT Mamba baseline. Here the hyperparameter setting for ReMamba is: $s=0$, $p=0.20$ and $\rho=0.05$. The results for DeciMamba2.8b also show no significant improvements.
5.4 Analyses and Discussions
5.4.1 Ablation Study
@@
Figure 4: Average scores on LongBench varying max length from 2k to 9k. The “Pre” means pretrained model while “SFT” means finetuned model. The performance of Llama-3b (SFT) and Llama-3b (Pre) is for reference, using the max length of 6k.
Figure 5: Average scores on L-Eval varying max length from 2k to 9k. The performance of Llama-3b (SFT) and Llama-3b (Pre) is for reference, using the max length of 6k.
Figure 6: Memory consumption comparisons during inference are conducted with a 6144 input sequence and a 1024 output sequence, using a batch size of 1. The experiments are done on an A100 80GB GPU.
5.4.2 Varying Length
To complement our main results, which employ a maximum sequence length of 6k tokens to align with training settings, we further evaluate the model performance at varying input lengths ranging from 2k to 9k tokens. This evaluation is conducted using the LongBench and L-Eval benchmarks. As depicted in Figure 4, our ReMamba consistently outperforms the baseline Mamba model across all tested context lengths on LongBench. Notably, the performance gap between our model and the baseline widens as the context length increases. Furthermore, our model extends the efficient context length (the length at which greatest performance is observed) to 6k tokens, compared to 4k tokens for the finetuned Mamba baseline. In Figure 5, we observe performance improvements across all context lengths for our model on L-Eval.
Model 2WikiMQA GovReport HotpotQA LCC MultiNews MultiQA PassCount PassRetrie. Qasper RepoBench SAMSum TREC TriviaQA Average Mamba2 (Pre) 2.18 4.76 1.54 23.46 7.71 2.88 0.60 1.47 1.17 14.97 2.07 8.33 10.60 6.29 Mamba2 (SFT) 13.73 19.97 15.05 41.78 19.52 25.20 0.67 5.67 13.44 38.79 33.95 49.67 43.54 24.69 ReMamba2 (SFT) 18.90 19.03 18.09 51.15 17.68 25.82 3.33 5.00 14.84 43.99 23.55 44.00 56.90 26.33
5.4.3 Speed Performance and Memory Expense
Our model introduces a single additional forward pass during inference, resulting in small constant memory consumption. We visualize the memory consumption during inference with a 6k input sequence and a 1k output sequence, using a batch size of 1 in Figure 6. The device we use is an A100 80GB GPU. We observe that the encoding process of Llama consumes a substantial amount of memory and the KV cache increases gradually during the decoding process, whereas ReMamba incurs only a small additional memory cost. After the encoding process, ReMamba’s memory consumption stabilizes at a constant level, which is moderately higher than Mamba’s, corresponding to the additional parameters introduced to support the selection mechanism.
To evaluate the speed performance, we varys the input sequence length from 1k to 8k tokens while fixing the output length at 1k tokens. For all the experiments, we use a batch size of 1 and measure the speed on an NVIDIA A100 80GB GPU. We compare the performance of ReMamba, Mamba, and the vanilla transformer model (Llama-3b), as illustrated in Figure 7. The speed metric is given in tokens per second. Our experiments indicate that ReMamba operates at speeds comparable to the original baseline, maintaining a significant speed advantage over traditional transformers.
5.4.4 Robustness varying choices of hyperparamters
The aforementioned results were obtained using the hyperparameter settings $s=0$, $p=0.18$, and $\rho=0.009$, which demonstrates relatively superior performance. In Figure 8, we also show the stability of our model by varying the hyperparameters $p$ and $\rho$. For these experiments, the hyperparameter $s$ is fixed at 0.
@@ 5.4.4 Robustness varying choices of hyperparamters
5.4.5 Generalizing to Mamba2
We also evaluated the applicability of our approach to Mamba2. As shown in Table 3, ReMamba achieves an average improvement of 1.6 points on LongBench. Further details are available in Appendix A.2.
6 Conclusions
Limitations
Although ReMamba improves the long-context performance of Mamba, it is unlikely that Mamba can outperform transformers as the context length increases, due to its fixed-size state space. Additionally, ReMamba primarily relieve the loss of long-context information in Mamba through selective compression. A more promising approach would be to directly modify the state space update mechanism.
Appendix A Appendix
Model dataset epoch optimizer learning scheduler learning rate warm up steps LoRA rank
Llama-3b (SFT) LongOrca 1 AdamW linear 2e-5 0 32
DeciMamba (SFT) LongOrca 1 AdamW linear 2e-5 0 32
Mamba (SFT) LongOrca 1 AdamW linear 2e-5 0 32
ReMamba (SFT) LongOrca 1 AdamW cosine 2e-5 0 32
Mamba2 (SFT) LongOrca 1 AdamW cosine 2e-5 0 32
ReMamba2 (SFT) LongOrca 1 AdamW cosine 2e-5 0 32
Table 4: Model training details
@@ Appendix A Appendix
Table 5: Performance of different model variants on LongBench. Parameters $s$, $p$, and $\rho$ represent the relative start position, relative length to compress, and compression ratio, respectively. In this context, the “ReMamba” model type constitutes our optimal model. The “middle” type corresponds to the model variant where $s$ is non-zero. The “special” model variant compresses the entire prompt using $\rho=0.009$ and subsequently appends the compressed hidden states to the end of the original prompt in the second stage.
A.1 Training Details
During training, the ReMamba hyperparameter $s$ is fixed at 0. The hyperparameter $p$ is randomly sampled from the interval [0.1, 0.3], while $\rho$ is randomly sampled from the interval [0.05, 0.2].
Most of DeciMamba’s experiments are conducted on Mamba130m. The only configuration for DeciMamba2.8 provided in the original paper is decimating_layers = 22, decimation_max_p_L_base = 4000, used for language modelling tasks. Here we change the decimation_max_p_L_base = 6000. For all other hyperparameters, we use the default settings: decimation_min_seq_len = 20, decimation_beta = 0.5, and decimation_type = “max_p“. The linear positional interpolation configurations for Llama-3B are: max_position_embeddings = 6144 and factor = 3. The data construction process for Llama-3b is identical to that of Mamba. We finetune all the models for 1 epoch using DeepSpeed Zero Stage 3 on 8 A100-80GB GPUs. Other training details can be found in the Table 4.
A.2 Mamba2 Details
The hyperparameters here are: $s=0$, $p=0.25$ and $\rho=0.05$. The max length is still 6k. It is noteworthy that Mamba2 exhibits nearly no performance improvement over Mamba on LongBench, suggesting potential limitations within the Mamba model series.
A.3 Why compress from the start